OpenStudy (anonymous):

Given f(x) = x^3 and g(x) = 2x + 1 where f,g: Q->Q, show that gof is a bijection. I plug in f into g, and idk what to do from there. Do I need to show that it's 1-1 and onto?

OpenStudy (rock_mit182):



OpenStudy (rock_mit182):

is this linear algebra ?

OpenStudy (anonymous):

Discrete math

OpenStudy (amistre64):

since f and g both have inverses, they are 1 to 1 since they are both odd degrees, they have the same domain/range as well, so onto. not sure how rigorous that is tho

OpenStudy (amistre64):

hmm, show that g(f) is also 1-1 and onto .... g(f) = 2x^3 + 1 is just a modified f ...

OpenStudy (amistre64):

2(a/b)^3 + 1 = 2(m/n)^3 + 1 2(a/b)^3 = 2(m/n)^3 (a/b)^3 = (m/n)^3 now we are left to showing that f is 1-1 and onto .... maybe